Consider a generic weak acid solution, HX(l) of pH 2.122 prepared by dissolving 8.42 g of solid HX(s) sample in 35.00 mL of water. What is the molar mass of this acid? Ka for HX is 4.0 × 10–5.
1
Expert's answer
2020-12-10T07:50:01-0500

pH= 2.122.

[H+] = 0.00755092227665 M.

Ka=a^2 x C.

a=√(Ka/C) = √(4.0 x 10-5/0.00755092227665) = 0,0727830072

0.00755092227665/0,0727830072 =0,0727830072 M.

Cm=n/V. n= Cm x V= 0,0727830072 x 0.035 = 0,00363109866 moles.

Mr=m/n=8.42/0,00363109866=2 318,8574 / 100 = 23,188574